OpenStudy (emilyjones284):

what are the points of discontinuity for y= x+1 over x^2-4?

OpenStudy (campbell_st):

you have vertical asymptotes when the denominator is zero \[y = \frac{x +1}{(x -2)(x + 2)}\] so you need to solve x - 2 = 0 and x + 2 = 0 there is a horizontal asymptote at y = 0 hope this helps

OpenStudy (emilyjones284):

thank you! (: